Let G(x)=(2+x^(1/3))^8 and suppose that we write G(x)=f(g(h(x))) with h(x)=x^(1/3), g(x)=A+x and f(x)=x^B
A=_____________ B=______________

Answer by stanbon(75887)

Let G(x)=(2+x^(1/3))^8 and suppose that we write G(x)=f(g(h(x))) with h(x)=x^(1/3), g(x)=A+x and f(x)=x^B
A=2_____________ B=8______________
----
1st machine is h(x) = x^(1/3)
---
2nd machine is g(x) = A+x =
means g(x^(1/3)) = x^1/3 + A
There you can see that A must be 2.
-----
3rd machine is the last machine.
It raised (2+x^(1/3)) to the 8th power.
So B = 8
================
Cheers,
Stan H.